OpenStudy (anonymous):

If f(x) = −2x − 4 and g(x) = x + 4, find (f/g)(−8).

OpenStudy (rulnick):

(-2(-8)-4)/(-8+4) = 12/(-4) = -3

OpenStudy (anonymous):